Find the hcf of 65 and 117 and express in the form 65m + 117n.

Grade:10

2 Answers

Yash rawat
6 years ago
Hcf of 65 and 117 is 13 By using euclid`s division lemma or prime factorizationTherefore 13=65m-11765m=117+1365m=130M=130÷65M=2
Kochouseph
6 years ago
65=13*5117=13*3*3Hcf of 65 and 117=1313=65m*117nHence the hcf of 65 and117 is 13 it is proved by Euclid division lemma